# Networked Communication for Mean-Field Games with Function Approximation and Empirical Mean-Field Estimation

### v2

@@ C.1 Games
This is the inverse of the ‘exploration’ game in Laurière et al. [2022b], where in our case agents are encouraged to gather together by the reward function $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=$ log$(\hat{\mu}_{t}(s^{i}_{t}))$. That is, agent $i$ receives a reward that is logarithmically proportional to the fraction of the population that is co-located with it at time $t$. We give the population no indication where they should cluster, agreeing this themselves over time.
Agree on a single target.
Unlike in the above ‘cluster’ game, the agents are given options of locations at which to gather, and they must reach consensus among themselves. If the agents are co-located with one of a number of specified targets $\phi\in\Phi$ (in our experiments we place one target in each of the four corners of the grid), and other agents are also at that target, they get a reward proportional to the fraction of the population found there; otherwise they receive a penalty of -1. In other words, the agents must coordinate on which of a number of mutually beneficial points will be their single gathering place. Define the magnitude of the distances between $x,y$ at $t$ as $dist_{t}(x,y)$. The reward function is given by $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=r_{targ}(r_{collab}(\hat{\mu}_{t}(s^{i}_{t})))$, where
$r_{targ}(x)=\begin{cases}x\quad&\text{if}\,\exists\phi\in\Phi\text{ s.t. dist}_{t}(s^{i}_{t},\phi)=0\\ -1\quad&\text{otherwise,}\\ \end{cases}$
$r_{collab}(x)=\begin{cases}x\quad&\text{if}\,\hat{\mu}_{t}(s^{i}_{t})>1/N\\ -1\quad&\text{otherwise.}\\ \end{cases}$
@@
Define the magnitude of the horizontal and vertical distances between $x,y$ at $t$ as $dist^{h}_{t}(x,y)$ and $dist^{v}_{t}(x,y)$ respectively. The state $s^{i}_{t}$ now consists of agent $i$’s position $x^{i}_{t}$ and a ‘shark’s’ position $\phi_{t}$. At each time step, the shark steps towards the most populated grid point according to the empirical mean-field distribution i.e. $x^{*}_{t}=\arg\max_{x\in\mathcal{S}}\hat{\mu}_{t}(x)$. A horizontal step is taken if $dist^{h}_{t}(\phi_{t},x^{*}_{t})$geq $dist^{v}_{t}(\phi_{t},x^{*}_{t})$, otherwise a vertical step is taken. As well as featuring a non-stationary distribution, we add ‘common noise’ to the environment, with the shark in a random direction with probability 0.01. Such noise that affects the local states of all agents in the same way, making the evolution of the distribution stochastic, makes population-independent policies sub-optimal Laurière et al. [2022a]. Agents are rewarded more for being further from the shark, and also for clustering with other agents. The reward function is given by
@@
This is similar to the ‘exploration’ tasks in Laurière et al. [2022b], Wu et al. [2024] and other MFG works. In our version agents are rewarded for being located in more sparsely populated areas but only if they are stationary. The reward function is given by $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=r_{stationary}(-\hat{\mu}_{t}(s^{i}_{t}))$, where
$r_{stationary}(x)=\begin{cases}x\quad&\text{if}\,a^{i}_{t}\text{ is `remain stationary'}\\ -1\quad&\text{otherwise.}\\ \end{cases}$
C.2 Experimental metrics
To give as informative results as possible about both performance and proximity to the MFNE, we provide two metrics for each experiment. Both metrics are plotted with mean and standard deviation, computed over the ten trials (each with a random seed) of the system evolution in each setting.
C.2.1 Exploitability
Works on MFGs most commonly use the exploitability metric to evaluate how close a given policy $\pi$ is to a NE policy $\pi^{*}$ Laurière et al. [2022b]; Perrin et al. [2020]; Laurière et al. [2022a]; Algumaei et al. [2023]; Pérolat et al. [2022]; Benjamin and Abate [2023]; Wu et al. [2024]. The metric usually assumes that all agents are following the same policy $\pi$, and quantifies how much an agent can benefit by deviating from $\pi$ by measuring the difference between the return given by $\pi$ and that of a $BR$ policy with respect to the distribution generated by $\pi$:
Definition 9 (Exploitability of $\pi$).
The exploitability ${E}x$ of policy $\pi$ is given by:
@@ C.2.1 Exploitability
If $\pi$ has a large exploitability then an agent can significantly improve its return by deviating from $\pi$, meaning that $\pi$ is far from $\pi^{*}$, whereas an exploitability of 0 implies that $\pi=\pi^{*}$. Prior works conducting empirical testing have generally focused on the centralised setting, so this classical definition, as well as most evaluations, only consider exploitability when all agents are following a single policy $\pi_{k}$. However, Benjamin and Abate [2023] notes that purely independent agents, as well as networked agents, may have divergent policies $\pi^{i}_{k}\neq\pi^{j}_{k}$ $\forall i,k\in 1,\dots,N$, as in our own setting. We therefore are interested in the ‘exploitability’ of the population’s joint policy $\boldsymbol{\pi}$ := ($\pi^{1},\dots,\pi^{N}$) $\in\Pi^{N}$.
@@
Since we do not have access to the exact $BR$ policy as in some related works Laurière et al. [2022b]; Wu et al. [2024], we must instead approximate the exploitability, similarly to Perrin et al. [2021]; Benjamin and Abate [2023]. We freeze the policy of all agents apart from a deviating agent, for which we store its current policy and then conduct 50 $k$ loops of policy improvement. To approximate the expectations in Def. 9, we take the best return of the deviating agent across 10 additional $k$ loops, as well as the mean of all the other agents’ returns across these same 10 loops. (While the policies of all non-deviating agents is $\pi_{k}$ in the centralised case, if the non-deviating agents do not share a single policy, then this method is in fact approximating the exploitability of their joint policy $\boldsymbol{\pi}^{-d}_{k}$, where $d$ is the deviating agent.) We then revert the agent back to its stored policy, before learning continues for all agents as per the main algorithm. Due to the expensive computations required for this metric, we evaluate it every second $k$ iteration of the main algorithm for Figs. 1, 2, 3, 4 and 5, and every fourth iteration for the population-dependent experiments.
The exploitability metric has a number of limitations in our setting. Our approximation takes place via MOMD policy improvement steps (as in the main algorithm) for an independent, deviating agent while the policies of the rest of the population are frozen. As such, the quality of our approximation is limited by the number of policy improvement/expectation rounds, which must be restricted for the sake of running speed of the experiments. Moreover, since one of the findings of our paper is that networked agents can improve their policies faster than independent or centralised agents, especially when non-linear function approximation is used, it is arguably unsurprising that approximating the $BR$ by an independently deviating agent sometimes gives an unclear and noisy metric. This includes the exploitability going below zero, which should not be possible if the policies and distributions are computed exactly.
@@
Moreover, in coordination games (the setting for all tasks apart from the ‘disperse’ game), agents benefit by following the same behaviour as others, and so a deviating agent generally stands to gain less from a $BR$ policy than it might in the non-coordination games on which many other works focus. For example, the return of a best-responding agent in the ‘push object’ game still depends on the extent to which other agents coordinate on which direction in which to push the box, meaning it cannot significantly increase its return by deviating. This means that the downward trajectory of the exploitability metric is less clear in our plots than in other works. This is likely why the approximated exploitability gets lower in the non-coordination ‘disperse’ task in Fig. 5 than in the other tasks. Given the limitations presented by approximating exploitability, we also provide the second metric to indicate the progress of learning.
C.2.2 Average discounted return
@@
We record the average discounted return of the agents’ policies $\pi_{k}^{i}$ during the $M$ iterations - this allows us to observe that settings that converge to similar exploitability values may not have similar average agent returns, suggesting that some algorithms are better than others at finding not just NE, but preferable NE. See for example Figs. 3 and 1, where the networked agents converge to similar exploitability as the independent and centralised agents, but receive higher average returns.
C.3 Hyperparameters
Figure 3: ‘Target agreement’ task, population-independent policies, 50x50 grid.
Figure 4: ‘Cluster’ task, population-independent policies, 50x50 grid.
Figure 5: ‘Disperse’ task, population-independent policies, 100x100 grid.
@@ C.3 Hyperparameters
In our experiments we generally want to demonstrate that our communication-based algorithms outperform the centralised and independent architectures by allowing policies that are estimated to be better performing to proliferate through the population, such that convergence occurs within fewer iterations and computationally faster, even when the Q-function is poorly approximated and/or the mean-field is poorly estimated, as is likely to be the case in real-world scenarios. Moreover we want to show that there is a benefit even to a small amount of communication, so that communication rounds themselves do not excessively add to time complexity. As such, we generally select hyperparameters at the lowest end of those we tested during development, to show that our algorithms are particularly successful given what might otherwise be considered ‘undesirable’ hyperparameter choices.
@@
Gridsize 10x10 / 50x50 / 100x100 Experiments with population-dependent policies are run on the 10x10 grid (Figs. 6, 7, 8 and 9), while experiments on large state spaces are run on 50x50 and 100x100 grids (Figs. 1, 2, 3, 4 and 5).
Population 500 We chose 500 for our demonstrations to show that our algorithm can handle large populations, indeed often larger than those demonstrated in other mean-field works, especially for grid-world environments, while also being feasible to simulate wrt. time and computation constraints Yongacoglu et al. [2022]; Cui and Koeppl [2021]; Cui et al. [2023]; Guo et al. [2023]; Subramanian and Mahajan [2019]; Yang et al. [2018]; Ganapathi Subramanian et al. [2021, 2020]; Subramanian et al. [2022]; Wu et al. [2024]; Benjamin and Abate [2023].
Number of neurons in input layer cf. comment The agent’s position is represented by two concatenated one-hot vectors indicating the agent’s row and column. An additional two such vectors are added for the shark’s/object’s position in the ‘evade’ and ’push object’ tasks. For population-dependent policies, the mean-field distribution is a flattened vector of the same size as the grid. As such, the input size in the ‘evade’ and ’push object’ tasks is $[(4\times\mathrm{dimension})+(\mathrm{dimension}^{2})]$; in the other settings it is $[2\times\mathrm{dimension}]$.
@@
$C_{p}$ 1 As in Benjamin and Abate [2023], we choose this value to show the convergence benefits brought by even a single communication round, even in networks that may have limited connectivity; higher choices are likely to have even better performance.
@@
$\tau_{q}$ 0.03 We tested $\tau_{q}$ in {0.01,0.02,0.03,0.04,0.05}, as well as linearly decreasing $\tau_{q}$ from 0.05 $\rightarrow$ 0 as $k$ increases. However, only 0.03 gave stable increase in return. Note that this is the value also chosen in Vieillard et al. [2020].
@@
$cl$1 We use the same value as in Vieillard et al. [2020].
@@
$\nu$ $L-1$ We tested $\nu$ in $\{1,4,20,L-1\}$. We found that in our setting, updating $\theta^{\prime}\leftarrow\theta$ once per $k$ iteration s.t. $\theta^{\prime}_{k+1,l}=\theta_{k,l}$ $\forall l$ gave sufficient learning that was similar to the other potential choices of $\nu$, so we do this for simplicity, rather than arbitrarily choosing a frequency to update $\theta^{\prime}$ during each $k$ loop. Setting the target to be the policy from the previous iteration is similar to the method in Laurière et al. [2022b]. Whilst Wu et al. [2024] updates the target within the $L$ loops for stability, we do not find this to be a problem in our experiments.
Optimiser Adam As in Vieillard et al. [2020], we use the Adam optimiser with initial learning rate 0.01.
@@
$\tau^{comm}_{k}$ cf. comment $\tau^{comm}_{k}$ increases linearly from 0.001 to 1 across the $K$ iterations. This is a simplification of the annealing scheme used in Benjamin and Abate [2023]. Further optimising the annealing process may lead to better results.
Table 1: Hyperparameters
C.4 Additional experiments
C.4.1 Additional experiments on large grids
We provide additional experiments on large grids in Figs. 3, 4 and 5.
In the ‘target agreement’ task in Fig. 3, the networked agents generally have lower exploitability than both centralised and independent agents, and significantly outperform the other architectures in terms of average return. As before, the margin by which the networked agents can outperform the centralised agents is much greater than in Benjamin and Abate [2023], showing that the benefits of the communication scheme are even greater in non-tabular settings.
In the ‘cluster’ task in Fig. 4, the networked agents obtain significantly higher return than the independent agents. While centralised agents have the lowest exploitability, networked agents of almost all communication radii outperform them in terms of average return, indicating that the communication scheme helps populations reach better performing equilibria.
@@ C.4.1 Additional experiments on large grids
In the ‘disperse’ task in Fig. 5, networked agents significantly outperform independent and centralised agents in terms of average return. They also outperform centralised agents in terms of exploitability, and significantly outperform independent agents in terms of exploitability. The fact that this happens in this non-coordination, competitive game shows that agents do have an incentive to communicate with each other even if they are self-interested.
C.4.2 Population-dependent policies in complex environments
Figure 6: ‘Push object’ task, population-dependent policies with estimated mean-field distribution, 10x10 grid.
Figure 7: ‘Evade’ task, population-dependent policies with estimated mean-field distribution, 10x10 grid.
Figure 8: ‘Push object’ task, population-dependent policies with global observability of mean field, 10x10 grid.
Figure 9: ‘Evade’ task, population-dependent policies with global observability of mean field, 10x10 grid.
We also showcase the ability of our algorithm to handle two more complex tasks, using population-dependent policies and estimated mean-field observations.
Figs. 6 and 7, where agents estimate the mean-field distribution via Alg. 3, differ minimally from Figs. 8 and 9, where agents directly receive the global mean-field distribution. This shows that our estimation algorithm allows agents to appropriately estimate the distribution, even with only one round of communication for agents to help each other improve their local counts. Only in the ‘push object’ task in Fig. 6, and there only with the smaller broadcast radii, do agents slightly underperform the returns of agents in the global observability case in Fig. 8, as is reasonable.
For the reasons given in Sup. Mat. C.2.1, the exploitability metric gives limited information in the ‘push object’ task in Fig. 6. In the ‘evade’ task in Fig. 7, exploitability suggests that centralised learners outperform the other cases. However, all of the networked cases significantly outperform the independent learners in terms of the average return to which they converge in both tasks. In the ‘push object’ task networked learners also significantly outperform centralised learners in all but the case with the smallest broad communication radius, while in the ‘evade’ task all networked cases perform similarly to the centralised case. Recall though that in the real world a centralised architecture is a strong assumption, a computational bottleneck and single point of failure.
Appendix D Related work
MFGs are a quickly growing research area, so we only discuss the works most closely related to this present work, and instead refer the reader to Benjamin and Abate [2023] for detailed discussion around the setting of networked communication for MFGs, and to Laurière et al. [2022a] for a broader survey of MFGs. Our work is most closely related to Benjamin and Abate [2023], which introduced networked communication to the infinite-horizon MFG setting. However, this work focuses only on tabular settings rather than using function approximation as in ours, and only addresses population-independent policies.
Laurière et al. [2022b] uses Munchausen Online Mirror Descent (MOMD), similar to our method for learning with neural networks, but there are numerous differences to our setting: most relevantly, they study a finite-horizon episodic setting, where the mean-field distribution is updated in an exact way and an oracle supplies a centralised learner with rewards and transitions for it to learn a population-independent policy. Wu et al. [2024] uses MOMD to learn population-dependent policies, albeit also with a centralised method that exactly updates the mean-field distribution in a finite-horizon episodic setting. Perrin et al. [2022] learns population-dependent policies with function approximation in infinite-horizon settings like our own, but does so in a centralised, two-timescale manner without using the empirical mean-field distribution. Zhang et al. [2024] also uses function approximation along a non-episodic path, but involves a generic central agent learning an estimate of the mean field rather than using an empirical population. Approaches that directly update an estimate of the mean field must be able to generate rewards from this arbitrary mean field, even if they otherwise claim to be oracle-free. They are thus inherently centralised algorithms and rely on strong assumptions that may not apply in real-world problems. Conversely, we are interested in practical convergence in online, deployed settings, where the reward is computed from the empirical finite population.
Yongacoglu et al. [2022] addresses decentralised learning from a continuous, non-episodic run of the empirical system using either full or compressed information about the mean-field distribution, but agents are assumed to receive this information directly, rather than estimating it locally as in the algorithm we now present. They also do not consider function approximation or inter-agent communication in their algorithms. In the closely related but distinct area of Mean-Field RL, Ganapathi Subramanian et al. [2021] does estimate the empirical mean-field distribution from the local neighbourhood, however agents are seeking to estimate the mean action rather than the mean-field distribution over states as in our MFG setting. Their agents also do not have access to a communication network by which they can improve their estimates.
Appendix E Limitations and future work

### v3

@@ A.1. Technical description of tasks
[…]
Target agreement.
Unlike in the above ‘cluster’ game, the agents are given options of locations at which to gather, and they must reach consensus among themselves. If the agents are co-located with one of a number of specified targets $\phi\in\Phi$ (in our experiments we place one target in each of the four corners of the grid), and other agents are also at that target, they get a reward proportional to the fraction of the population found there; otherwise they receive a penalty of -1. In other words, the agents must coordinate on which of a number of mutually beneficial points will be their single gathering place to maximise their individual rewards. Define the magnitude of the distances between $x,y$ at $t$ as $dist_{t}(x,y)$. The reward function is given by $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=r_{targ}(r_{collab}(\hat{\mu}_{t}(s^{i}_{t})))$, where
[…]
@@
This is a similar idea to the task found in liu2025achievement. Define the magnitude of the horizontal and vertical distances between $x,y$ at $t$ as $dist^{h}_{t}(x,y)$ and $dist^{v}_{t}(x,y)$ respectively. The state $s^{i}_{t}$ now consists of agent $i$’s position $x^{i}_{t}$ and a ‘shark’s’ position $\phi_{t}$. At each time step, the shark steps towards the most populated grid point according to the empirical mean-field distribution i.e. $x^{*}_{t}=\arg\max_{x\in\mathcal{S}}\hat{\mu}_{t}(x)$. A horizontal step is taken if $dist^{h}_{t}(\phi_{t},x^{*}_{t})$ $\geq$ $dist^{v}_{t}(\phi_{t},x^{*}_{t})$, otherwise a vertical step is taken. As well as featuring a non-stationary distribution, we add ‘common noise’ to the environment, with the shark moving in a random direction with probability 0.01. Such noise that affects the local states of all agents in the same way, making the evolution of the distribution stochastic, makes population-independent policies sub-optimal (survey_learningMFGs). Agents are rewarded more for being further from the shark, and also for clustering with other agents. The reward function is given by
@@
The above tasks are all coordination tasks, in that agents receive higher returns by aligning their policies and hence have an incentive to communicate their policy parameters even though the MFG framework is non-cooperative. Our fifth game is a non-coordination task: the reward function is not designed to give higher returns for more aligned policies. We include this game to demonstrate that even in such non-cooperative scenarios our networked architecture receives higher returns than both independent and central-agent alternatives, such that decentralised selfish agents may still have incentive to communicate. The fifth game is:
@@
This is similar to the ‘exploration’ tasks in scalable_deep, wu2024populationaware and other MFG works. In our version agents are rewarded for being located in more sparsely populated areas, but only if they are stationary. The reward function is given by $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=r_{stationary}(-\hat{\mu}_{t}(s^{i}_{t}))$, where
[…]
A.2. Experimental metrics
To give as informative results as possible about both performance and proximity to the MFNE, we provide two metrics for each experiment. Both metrics are plotted with mean and standard deviation, computed over the ten trials (each with a random seed) of the system run in each setting.
A.2.1. Exploitability
Works on MFGs most commonly use the exploitability metric to evaluate how close a given policy $\pi$ is to a NE policy $\pi^{*}$ (scalable_deep; continuous_fictitious_play; survey_learningMFGs; algumaei2023regularization; scaling_MFG; benjamin2024networked; wu2024populationaware). The metric usually assumes that all agents are following the same policy $\pi$, and quantifies how much an agent can benefit by deviating from $\pi$ by measuring the difference between the return given by $\pi$ and that of a $BR$ policy with respect to the distribution generated by $\pi$:
Definition A.1 (Exploitability of $\pi$).
The exploitability Ex{E}x of policy $\pi$ is given by:
@@ A.2.1. Exploitability
If $\pi$ has a large exploitability then an agent can significantly improve its return by deviating from $\pi$, meaning that $\pi$ is far from $\pi^{*}$, whereas an exploitability of 0 implies that $\pi=\pi^{*}$. Prior works conducting empirical testing have generally focused on the central-agent setting, so this classical definition, as well as most evaluations, only consider exploitability when all agents are following a single policy $\pi_{k}$. However, benjamin2024networked notes that purely independent agents, as well as networked agents, may have divergent policies $\pi^{i}_{k}\neq\pi^{j}_{k}$ $\forall i,k\in 1,\dots,N$, as in our own setting. We therefore are interested in the ‘exploitability’ of the population’s joint policy $\boldsymbol{\pi}$ := ($\pi^{1},\dots,\pi^{N}$) $\in\Pi^{N}$.
@@
Since we do not have access to the exact $BR$ policy as in some related works (scalable_deep; wu2024populationaware), we must instead approximate the exploitability, similarly to flock; benjamin2024networked. We freeze the policy of all agents apart from a deviating agent, for which we store its current policy and then conduct 50 $k$ loops of policy improvement. To approximate the expectations in Def. A.1, we take the best return of the deviating agent across 10 additional $k$ loops, as well as the mean of all the other agents’ returns across these same 10 loops. (While the policies of all non-deviating agents is $\pi_{k}$ in the central-agent case, if the non-deviating agents do not share a single policy, then this method is in fact approximating the exploitability of their joint policy $\boldsymbol{\pi}^{-d}_{k}$, where $d$ is the deviating agent.) We then revert the agent back to its stored policy, before learning continues for all agents as per the main algorithm. Due to the expensive computations required for this metric, we evaluate it every second $k$ iteration of the main algorithm for Figs. 1, 2, 5, 7 and 6, and every fourth iteration for the population-dependent experiments.
@@
The exploitability metric has a number of limitations in our setting. In coordination games (the setting for all tasks apart from the ‘disperse’ game), agents benefit by following the same behaviour as others, and so a deviating agent generally stands to gain less from switching to a $BR$ policy than it might in the non-coordination games on which many other works focus. For example, the return of a best-responding agent in the ‘cluster’ task still depends on the extent to which other agents coordinate on where to cluster, meaning it cannot significantly increase its return by deviating from a badly clustering policy. This means that the downward trajectory of the exploitability metric is less clear in our plots than in other works that do not focus on coordination games. This is likely why the difference between the approximated exploitability of the independent agents and the other populations is clearer in the non-coordination ‘disperse’ task in Fig. 6 than in the other tasks.
Moreover, our approximation takes place via MOMD policy improvement steps (as in the main algorithm) for an independent, deviating agent while the policies of the rest of the population are frozen. As such, the quality of our approximation is limited by the number of policy-improvement/expectation-estimation rounds, which must be restricted for the sake of the running speed of the experiments. Moreover, since one of the findings of our paper is that independent-learning agents increase their returns significantly slower (if at all) than networked or central-agent populations, it is arguably unsurprising that approximating the $BR$ by an independently deviating agent sometimes gives an unclear and noisy metric. This includes the exploitability going below zero, which should not be possible if the policies and distributions are computed exactly. Given the limitations presented by approximating exploitability, we also provide a second metric to indicate the progress of learning, as in benjamin2024networked.
A.2.2. Average discounted return
@@
We record the average discounted return of the agents’ policies $\pi_{k}^{i}$ during the $M$ iterations. This allows us to compare how quickly agents are learning to increase their returns, even when exploitability gives us limited ability to distinguish between the desirability of the MFNEs to which populations converge. We can observe that settings that converge to similar exploitability values may not have similar average agent returns, suggesting that some algorithms are better than others at finding not just NE, but preferable NE (for reasons related to why networked populations can increase their returns faster as per Sec. 6). See for example Figs. 1 and 5, where the networked populations converge to similar exploitability as the independent and central-agent populations, but receive higher average returns.
A.3. Hyperparameters
@@
In our experiments we generally want to demonstrate that our communication-based algorithms outperform the central-agent and independent architectures by allowing policies that are estimated to be better performing to proliferate through the population, such that convergence occurs within fewer iterations and computationally faster, even when the Q-function is poorly approximated and/or the mean-field is poorly estimated, as is likely to be the case in real-world scenarios. Moreover we want to show that there is a benefit even to a small amount of communication, so that communication rounds themselves do not excessively add to time complexity. As such, we generally select hyperparameters at the lowest end of those we tested during development, to show that our algorithms are particularly successful given what might otherwise be considered ‘undesirable’ hyperparameter choices.
Table 1. Hyperparameters
@@
Gridsize $\times1010$/ $\times5050$/ $\times100100$ Experiments on large state spaces are run on $50\times 50$ and $100\times 100$ grids (Figs. 1-6), while experiments with population-dependent policies are run on the $10\times 10$ grid (Figs. 8(a), 9(a), 8(b) and 9(b)).
Population 500 We chose 500 for our demonstrations to show that our algorithm can handle large populations, indeed often larger than those demonstrated in other mean-field works, especially for grid-world environments, while also being feasible to simulate wrt. time and computation constraints (subjective_equilibria; approximately_entropy; cui2023multiagent; general_framework; rl_stationary; mfrl_yang; pomfrl; subramanian2022multi; DMFG; wu2024populationaware; benjamin2024networked).
Number of neurons in input layer cf. comment The agent’s position is represented by two concatenated one-hot vectors indicating the agent’s row and column. An additional two such vectors are added for the shark’s/object’s position in the ‘evade’ and ’push object’ tasks. For population-dependent policies, the mean-field distribution is a flattened vector of the same size as the grid. As such, the input size in the ‘evade’ and ’push object’ tasks is $(4\times\mathrm{dimension})+(\mathrm{dimension}^{2})]$; in the other settings it is $2\times\mathrm{dimension}]$.
@@
$C_{p}$ 1 As in benjamin2024networked, we choose this value to show the convergence benefits brought by even a single communication round, even in networks that may have limited connectivity; higher choices are likely to have even better performance.
@@
$\tau_{q}$ 0.03 We tested $\tau_{q}$ in {0.01,0.02,0.03,0.04,0.05}, as well as linearly decreasing $\tau_{q}$ from 0.05 $\rightarrow$ 0 as $k$ increases. However, only 0.03 gave stable increase in return. Note that this is the value also chosen in NEURIPS2020_2c6a0bae.
@@
clcl -1 We use the same value as in NEURIPS2020_2c6a0bae.
@@
$\nu$ L1L-1 We tested $\nu$ in $\{1,4,20,L-1\}$. We found that in our setting, updating $\theta^{\prime}\leftarrow\theta$ once per $k$ iteration s.t. $\theta^{\prime}_{k+1,l}=\theta_{k,l}$ $\forall l$ gave sufficient learning that was similar to the other potential choices of $\nu$, so we do this for simplicity, rather than arbitrarily choosing a frequency to update $\theta^{\prime}$ during each $k$ loop. Setting the target to be the policy from the previous iteration is similar to the method in scalable_deep. Whilst wu2024populationaware updates the target within the $L$ loops for stability, we do not find this to be a problem in our experiments.
Optimiser Adam As in NEURIPS2020_2c6a0bae, we use the Adam optimiser with initial learning rate 0.01.
@@
$\tau^{comm}_{k}$ cf. comment $\tau^{comm}_{k}$ increases linearly from 0.001 to 1 across the $K$ iterations. This is a simplification of the annealing scheme used in benjamin2024networked. Further optimising the annealing process may lead to better results.
A.4. Additional experiments and plots
@@
In the non-coordination ‘disperse’ task in Fig. 6, networked agents significantly outperform independent and central-agent populations in terms of average return. They also significantly outperform independent agents in terms of exploitability, and perhaps also central-agent populations though not significantly so. The fact that this happens in this non-coordination, competitive game shows that agents may have an incentive to communicate with each other even if they are self-interested. Indeed the agents learning independently do not appear to improve their returns at all, despite this being the paradigm that might be expected to perform best in a non-cooperative setting.
Appendix B Further discussion on the necessity of population-dependent policies
There are numerous reasons why agents may benefit from being able to respond to the current distribution. These include when:
  • There may in fact be any initial distribution (perrin2022generalization; wu2024populationaware; survey_learningMFGs).
  • The solved MFG has a non-stationary equilibrium, e.g. if there is a finite time horizon meaning that actions become more urgent as time passes such that policies (and therefore also the distribution) are non-stationary (scalable_deep; wu2024populationaware).
  • The distribution is subject to ‘common noise’. This is noise that affects the local states of all agents in the same way, making the evolution of the distribution stochastic. In such cases, even if the agent knows the policy used by all other agents, it cannot perfectly predict the evolution of the mean-field distribution, making population-independent policies suboptimal (carmona_common; cardaliaguet2015masterequationconvergenceproblem; survey_learningMFGs; li2024incomplete; becherer2024common).
  • The distribution deviates from the equilibrium for some other reason, such that agents need to be able to generalise their response to other (possibly previously unseen) distributions (survey_learningMFGs).
Appendix C Future work